# JWST COMPASS: Insights into the Systematic Noise Properties of NIRSpec/G395H From a Uniform Reanalysis of Seven Transmission Spectra

### v1

@@ 3.2 Wavelength Dependence
4 Analysis of Spectra
@@
4.1 Flat Spectral Models
@@
We begin our analysis by considering a model consisting of a flat line (i.e. a featureless spectrum) with an optional offset between the NRS1 and NRS2 portions of the spectra. We carry out a two generalized least squares fits for each spectrum: one with an offset and one with the offset fixed to zero. Table 7 contains the best-fit offset parameter, reduced chi-square for the fit with and without an offset, and the difference in the Bayesian information criteria between the two cases. Following Raftery1995, we find that the values indicate there is strong evidence for an offset for TOI-836 c, GJ 357 b, TOI-776 b, and TOI-776 c, while there is moderate evidence for the model without an offset for the remaining targets. The reduced chi-square values indicate that a flat spectrum model, either with or without an offset, is an adequate model for every spectrum. This result is consistent with previous analyses of each target that found that no putative features in the spectra rose to the level of a detection. We therefore proceed with our analysis under the assumption that there are no detectable molecular features in any of the spectra we present in this work.
| target | offset (ppm) | offset | no offset | (no offset - offset) |
| TOI-836 c | -34.23 | 1.87 | 1.15 | 21.74 |
| GJ 357 b | -61.12 | 2.63 | 0.98 | 71.62 |
| TOI-836 b | 6.14 | 1.31 | 1.28 | -3.58 |
| L 98-59 c | -5.93 | 1.03 | 1.02 | -3.94 |
| TOI-776 b | 45.82 | 3.08 | 1.24 | 57.29 |
| L 168-9 b | 0.38 | 1.38 | 1.39 | -4.61 |
| TOI-776 c | 20.57 | 1.31 | 0.86 | 15.36 |
4.2 Thermochemical Equilibrium Modeling
@@
Using our uniformly reduced multi-visit spectra, we carry out thermochemical equilibrium modeling following a methodology similar to that carried out previously for each target. The purpose of repeating this exercise is twofold: First, we wish to investigate the robustness of these lower limits to different choices made during the reduction and light curve fitting steps, and second, we wish to provide a set of lower limits derived from a uniform reduction and analysis. We begin by computing a grid of model atmospheres for each planet in terms of bulk metallicity and “opaque pressure level” (e.g., Wallack2024). The opaque pressure level is the pressure above which the atmosphere is clear, and below which all wavelengths of light are uniformly absorbed. In reality this could represent the top of a cloud deck, a haze layer, or a planetary surface. For our forward model we use Picaso (Batalha2019) with equilibrium chemical abundances computed by easyCHEM (Lei2024). While we use the full set of molecules output by easyCHEM to calculate the mean molecular weight of the atmosphere, we include only CO, H2O, CH4, PH3, CO2, NH3, H2S, VO, TiO, Na, K, Cs, Li, N2O, O2, O3, and Rb as absorbing species. We compute the equilibrium chemistry for C/O ratios of 0.5 Solar, 1.0 Solar, and 1.5 Solar (where Solar C/O=0.55; Asplund2009) in order to investigate the dependence of the resulting limits on C/O ratio. Table 5 lists the planetary parameters that were used to compute our forward models.
@@
where is the interpolated model spectrum based on our model grid and 3.8 is an arbitrary cutoff chosen to fall in the detector gap between NRS1 and NRS2. The prior on metallicity is uniform between 10 and 1000 Solar, and the prior on is uniform between -6 and -1. We do not place a prior on the offset parameters.
For each target we simulate 8 MCMC chains for 1,000,000 steps each, discarding the first 20,000 steps as burn-in (2% of the total simulation). This results in an effective sample size of at least 80,000 for each parameter based on an analysis of the autocorrelation length and parameters equal to 1.0 to three decimal places indicating that the simulations are well-converged, with both quantities computed using arviz (arviz). We see no evidence of multimodality in the posteriors.
@@
In order to control for the extent to which our methodology contributes to any disagreement between our results and those of previous COMPASS analyses, we also compute the same limits for metallicity and opaque pressure level using identical pipelines from the original works. As described in other works ((e.g. Wallack2024; Teske2025; Alderson2025)) we compute a similar model grid using PICASO and photochem (Wogan2025) to compute spectra at chemical equilibrium abundances. We choose two values of C/O (0.5Solar and 1Solar, where Solar=0.458 (2009LanB...4B..712L)) and 20 log-even spaced metallicities between M/H=[0,3]. We fit for an offset between NRS1 and NRS2 using ultranest ultranest. Then, we compute a reduced per data point of the NRS1/NRS2 offset corrected spectrum, which we can then be converted to a p-value and associated -confidence. The affect of clouds is added to the spectrum by including a thick () gray opacity source at a specified pressure level. Our grid includes 5 opaque pressure levels spanning 1-10-4 bars.
4.3 Composite spectra
@@
5 Atmospheric Composition Results
@@
5.1 Influence of C/O Ratio
For planets with equilibrium temperature below about 1000 K, changes in the C/O ratio have minimal impacts on the transmission spectrum (Madhusudhan2012). As most of the planets in our sample are in this regime, the results of our thermochemical equilibrium modeling are robust to changes in C/O, justifying the choice to assume C/O=1 Solar. However, TOI-836 b and L 168-9 b, with equilibrium temperatures of 870 K and 1000 K respectively, are warm enough to be influenced by the choice of C/O ratio. At these temperature the mixing ratios of H2O and CO2 can change substantially between C/O=1 Solar and C/O=1.5 Solar, as illustrated in Figure 15, which in turn drives changes in the posteriors for opaque pressure level versus metallicity.
@@ 5.1 Influence of C/O Ratio
In the case of L 168-9 b, for which we find a metallicity lower limit of Solar for the C/O=1 Solar case, the discrepancy is at least in part driven by the planet parameters used in our forward models. Alam2025 used the mass, radius, and equilibrium temperature from AstudilloDefru2020, while we use Hobson2024. When we substitute the AstudilloDefru2020 for our forward model grid computation, we find a metallicity lower limit of 335 Solar for the C/O=1 Solar case. This is significantly closer, but still much higher, than the limit of Solar found by Alam2025.
@@
While this sensitivity may partially explain the difference between the lower limits derived from the MCMC method and those derived from the -mapping method for L 98-59 c and L 168-9 b (the black and green vertical lines in Figure 17). However, it cannot explain the discrepancy between the lower limits derived in previous COMPASS studies and those found in this work since the two methods are in such close agreement for the other targets. A possible explanation is found in the fact that the change in the amplitude of transmission features with changes in metallicity grows smaller as metallicity increases. Our L 168-9 b and L 98-59 c spectra constrain the metallicity to be quite high, putting us in a regime where the feature amplitude, and therefore the of the models with respect to the data, are relatively insensitive to changes in metallicity. This means that large changes in the metallicity are necessary to effect small changes in the quality of the fit, making the metallicity lower limit sensitive to small changes in the spectrum between reductions as well as to the goodness of fit metric used (i.e. versus the log-likelihood).
5.2 Results of Composite Spectra Analysis
@@
To test this hypothesis we now present two composite spectra: One for the sub-Neptunes TOI-776 c and TOI-836 c, and the other for the five super-Earths (L 98-59 c, L 168-9 b, GJ 357 b, TOI-836 b, and TOI-776 b). Each composite spectra is produced by co-adding the group of spectra with error bars are obtained by summing the error bars for the individual spectra in quadrature. Offsets were not removed prior to co-adding the spectra, so the offset in the composite spectra are the sum of the individual offsets.
@@
6 Discussion
There are a range of possible explanations for the lack of transmission features for these seven COMPASS targets. The majority of the sample is made up of super-Earths, for which atmosphere detections have proven challenging (see, e.g. May2023; Cadieux2024; Moran2023, as well as previous COMPASS papers). Of the ten super-Earths observed by JWST in transmission in the near-infrared, only one claim of an atmospheric detection has been made, for L 98-59 b (Bello-Arufe2025). If these planets do possess atmospheres, they may be high mean molecular weight secondary atmospheres. Compositions with mean molecular weights greater than five to ten g/mol are permitted for all five super-Earths in our sample given the precision of the transmission spectra presented here. Future JWST investigations including the Rocky Worlds Director’s Discretionary Time program and the Charting the Cosmic Shoreline Guest Observer program (#7073, PI: Lustig-Yaeger) will help to answer lingering questions about whether these small worlds can harbor atmospheres, at least in the case of the M-dwarf host stars targeted by those programs.
The two sub-Neptunes in our sample, TOI-776 c and TOI-836 c, both have bulk densities that imply the presence of volatiles, and Helium escape has been detected for TOI-836 c, and Hydrogen escape has been observed for TOI-776 demonstrating that they must retain at least some of their initial H/He envelopes (Loyd2025; Zhang2025). TOI-836 c is a particularly puzzling case due to its proximity in mass-radius space to GJ 1214 b, TOI-421 b, and K2-18 b, all of which have atmosphere detections. However, TOI-836 c has an equilibrium temperature of 665K, which places it in a range where previous studies have suggested that hazes may play a role in attenuating spectral features (Brande2024; Crossfield2017). Indeed, Wallack2024 finds that haze production rates on par with those predicted for GJ 1214 b could produce a high-altitude haze for TOI-836 c.
@@ 6 Discussion
6.1 Comparison of COMPASS Targets to other Small Planets Observed with JWST
@@
The seven COMPASS targets studied in this work span from one to three Earth-radii in size. In this same radius range there are a total of six planets observed by JWST in the near-infrared which are thought to show spectral features. These are L 98-59 b (Bello-Arufe2025), GJ 9827 d (Piaulet2024), TOI-270 d (Benneke2024), K2-18 b (Madhusudhan2023), TOI-421 b (Davenport2025), and GJ 1214 b (Schlawin2024; Ohno2025). There are an additional five small planets observed by JWST in the near-infrared which do not have transmission features. These are GJ 3090 b (Ahrer2025), GJ 1132 b (May2023; Bennett2025), GJ 486 b (Moran2023), LHS 475 b (LustigYaeger2023), and LHS 1140 b (Cadieux2024).
@@
In addition to radius, these planets cover similar ranges in mass and equilibrium temperature to the COMPASS targets (see Figures 19 and 20). A possible trend emerges when we consider the distribution of planets with transmission features compared to those without: With a few exceptions (GJ 3090 b and TOI-836 c), the planets with transmission features tend to sit near or above the radius expected for a 50% steam atmosphere (see the dark blue line from Aguichine2021) while the planets without transmission features tend to sit below this line, closer to the pure silicate model from Zeng2016. A fuller understanding of the statistical significance of this trend and its implications for the exoplanet population would require an analysis that is beyond the scope of this work. However, this emerging trend suggests that atmospheres may be challenging to observe for denser worlds.
@@
We examine the transmission spectra that would be produced by the atmospheres of the six planets with features if they were placed around each of the COMPASS planets. We take the retrieved molecular abundances from the original studies, for which best-fit spectra are shown in the left panel of Figure 21, and use these to compute model spectra for the COMPASS targets using petitRADTRANS (Molliere2019). These models are shown in the right panel of Figure 21, where they are plotted over the observed spectra of each atmosphere. In Figure 22 we show the values between the observed planet spectra (columns) and the model atmosphere (rows). For GJ 1214 b we used the CO2 and CH4 abundances from Schlawin2024, who did not report a mean molecular weight. We found that in order to reproduce an approximate fit to the observations we had to assume a mean molecular weight of about 42 amu, which is roughly consistent with the reported abundances plus a background gas composed of N2. We note that the composition of GJ 1214 b’s atmosphere and haze properties are highly uncertain (see also Ohno2025 and Malsky2025).
@@
We find that the low mean molecular weight compositions of K2-18 b and TOI-421 b are inconsistent with the flat spectra of all seven COMPASS targets at greater than 3-. This is because their low mean molecular weight atmospheres produce large transmission features, which are easy to rule out for the COMPASS spectra. TOI-270 d has an intermediate mean molecular weight of =5.82 amu, which results in muted spectral features. Its composition is ruled out for the super-Earths (except for TOI-776 b), but remains barely plausible for TOI-836 c at and is fully consistent with the spectrum of TOI-776 c. The purported high mean molecular weight atmosphere of L 98-59 b is inconsistent with the flat spectrum of its sibling planet L 98-59 c, but is otherwise plausible for the remaining COMPASS targets. However, we note that there are uncertainties about the existence and composition of this atmosphere of L 98-59 b, with the observed JWST spectrum also being consistent with a bare rock (Bello-Arufe2025). The water-dominated atmosphere of GJ 9827 b is ruled out by the flat spectrum of the super-Earths L 168-9 b and disfavored for TOI-836 b and TOI-776 b, but is otherwise compatible with the remaining COMPASS planets. The very high mean molecular weight, hazy atmosphere of GJ 1214 b is flat enough to be plausible for all of the COMPASS planets.
@@
Summarizing these observations, we see that only the low mean molecular weight atmospheres observed for TOI-421 b and K2-18 b can be reliably ruled out for the entire sample. Higher mean molecular weight atmospheres are largely consistent with the flat spectra observed for these COMPASS targets, with several exceptions in the cases of L 98-59 c and L 168-9 b. In general, the flatness of the COMPASS spectra does not appear all that surprising in light of the compositional diversity and frequently high mean molecular weight atmospheres already observed among the collection of small planets seen by JWST.
6.2 Future Observations

### v2

@@ III.2 Wavelength Dependence
III.3 Nonlinearity and the Brighter-Fatter Effect
While the systematics model we have implemented is effective at mitigating correlated noise on short timescales, Figure 3 shows that the model does not account for all of the observed red noise. Residual red noise can also be seen in Figure 1 and the accompanying figure set. One possible culprit is the brighter-fatter effect (BFE). BFE is a phenomenon affecting infrared detectors wherein the accumulation of photoelectrons in a pixel near the center of an under-sampled PSF causes changes in the shape of the electric potential and repulses new incoming photoelectrons to adjacent pixels. In addition to causing the PSF to become fatter in brighter regions of the detector, BFE causes the charge as a function of groups to become nonlinear as the pixel gets closer to saturation (see Figure 14). The brighter-fatter effect has been studied for the MIRI Si:As IBC detectors (Argyriou et al., 2023; Gasman et al., 2024), for WFIRST’s H4RG-10 detectors (Hirata and Choi, 2020), and for Teledyne H4RG detectors (Plazas et al., 2018) and the Teledyne H2RG detectors used by JWST’s NIRISS and NIRSpec instruments (Goudfrooij et al., 2024; Alam et al., 2026). Alam et al. (2026) demonstrates that the brighter-fatter effect can acquire time-dependence in timeseries observations due to its dependence on the illumination of the detector. However the degree to which this effect impacts transit depth measurements is still understudied.
We investigate the presence of the brighter-fatter effect in our data using the observations from the second visit of TOI-776 b. Because these observations used seven groups, we are able to use the first several groups for which the accumulation of counts is relatively linear as a baseline to explore the nonlinearity of the last several groups. Starting with detector images that that have been destriped at the group level, we construct a median integration by taking the median detector image for each group across all integrations in the observation. Figure 16 shows the median accumulated counts as a function of group for two pixels on the NRS1 detector: one near the center of the trace (blue) and its neighboring pixel (red). The location of this central pixel is identified on the NRS1 detector in Figure 15 by the intersection of the horizontal and vertical dashed lines. To highlight the nonlinearity caused by the brighter-fatter effect we have fit a linear model to the first three groups and the examined the residuals between this model and all seven groups. These residuals are shown in the bottom panel of the figure. For the central pixel we see that the counts start out linear, but fall increasingly below the linear model as the integration proceeds. For its neighboring pixel the counts begin linear and then fall increasingly above the linear model. This is the expected behavior for the brighter-fatter effect: As the charge accumulates in the central pixel, photoelectrons that would otherwise fall on that pixel are diverted to neighboring pixels resulting in an excess of flux for the neighbors and a deficit for the central pixel which only gets worse as the integration proceeds.
We construct a diagnostic of the BFE for each pixel, denoted which approximates the second derivative of the counts () as a function of group number () at the location of the second to last group ( where is the number of groups for the observation). In Figure 15 we show how this value varies across the NRS1 and NRS2 detectors. We see that the approximated second-derivative is negative near the center of the trace in both NRS1 and NRS2, and then becomes positive at the edges of the trace, again as expected given our understanding of the brighter-fatter effect. Notably, the effect appears much stronger in NRS1 than NRS2.
We compute the difference between the maximum and minimum values of for detector row 16 in NRS1 and row 11 in NRS2 for the first visit of each target. We normalize this value by the median of , the average slope of counts between the first and second groups, taken over all pixels. This normalization is necessary if we wish to compare charge migration between different targets, because higher count rates from brighter stars will always result in more charge migration between groups. We choose to normalize by the slope at the beginning of the integration because the ramp is expected to be linear at this point, although the expectation of linearity at the beginning of the count accumulation ramp may break down for observations with very small group numbers. These values, which are plotted in Figure 16, indicates the extent to which the brighter-fatter effect impacts the linearity of the up-the-ramp samples for each observation. The first thing we note is that the nonlinearity varies more from observation to observation for NRS1 than NRS2. It also appears to be the case that the observations with seven groups experience less, and more consistent, charge migration than those observations with three or four groups. However, it is difficult to make precise statements about trends in charge migration with group number for such a small sample of observations.
While a correction step to mitigate the brighter-fatter effect has been implemented for NIRISS in the JWST calibration pipeline (Goudfrooij et al., 2024), there is as of yet no correction for NIRSpec. While a deeper exploration of the impacts of the brighter-fatter effect on transit depth measurements and ultimately the transmission spectrum itself is beyond the scope of this work, it is clear that the brighter-fatter effect impacts merits dedicated investigation in future work.
IV Analysis of Spectra
@@
IV.1 Flat Spectral Models
@@
We begin our analysis by considering a model consisting of a flat line (i.e. a featureless spectrum) with an optional offset between the NRS1 and NRS2 portions of the spectra. We carry out a two generalized least squares fits for each spectrum: one with an offset and one with the offset fixed to zero. Table 7 contains the best-fit offset parameter, reduced chi-square for the fit with and without an offset, and the difference in the Bayesian information criteria between the two cases. Following Raftery (1995), we find that the values indicate there is strong evidence for an offset for TOI-836 c, GJ 357 b, TOI-776 b, and TOI-776 c, while there is moderate evidence for the model without an offset for the remaining targets. The reduced chi-square values indicate that a flat spectrum model, either with or without an offset, is an adequate model for every spectrum. This result is consistent with previous analyses of each target that found that no putative features in the spectra rose to the level of a detection. We therefore proceed with our analysis under the assumption that there are no detectable molecular features in any of the spectra we present in this work.
| target | offset (ppm) | no offset | offset | (no offset - offset) |
| TOI-836 c | -46.335.99 | 1.87 | 1.10 | 21.96 |
| GJ 357 b | -66.275.75 | 2.63 | 0.96 | 72.26 |
| TOI-836 b | 4.184.98 | 1.31 | 1.26 | -3.94 |
| L 98-59 c | -0.544.93 | 1.03 | 0.99 | -2.70 |
| TOI-776 b | 43.074.58 | 3.08 | 1.23 | 57.63 |
| L 168-9 b | 14.164.32 | 1.38 | 1.27 | -2.70 |
| TOI-776 c | 19.264.19 | 1.31 | 0.86 | 15.67 |
IV.2 Thermochemical Equilibrium Modeling
@@
Using our uniformly reduced multi-visit spectra, we carry out thermochemical equilibrium modeling following a methodology similar to that carried out previously for each target. The purpose of repeating this exercise is twofold: First, we wish to investigate the robustness of these lower limits to different choices made during the reduction and light curve fitting steps, and second, we wish to provide a set of lower limits derived from a uniform reduction and analysis. We begin by computing a grid of model atmospheres for each planet in terms of bulk metallicity and “opaque pressure level” (e.g., Wallack et al., 2024). The opaque pressure level is the pressure above which the atmosphere is clear, and below which all wavelengths of light are uniformly absorbed. In reality this could represent the top of a cloud deck, a haze layer, or a planetary surface. For our forward model we use Picaso (Batalha et al., 2019) with equilibrium chemical abundances computed by easyCHEM (Lei and Mollière, 2024). While we use the full set of molecules output by easyCHEM to calculate the mean molecular weight of the atmosphere, we include only CO, H2O, CH4, PH3, CO2, NH3, H2S, VO, TiO, Na, and K as absorbing species, with opacities from the Zenodo v2 database of Resampled Opacities (Batalha et al., 2025). We compute the equilibrium chemistry for C/O ratios of 0.5 Solar, 1.0 Solar, and 1.5 Solar (where Solar C/O=0.55; Asplund et al., 2009) in order to investigate the dependence of the resulting limits on C/O ratio. Table 4 lists the planetary parameters that were used to compute our forward models. The forward models are computed at and then binned to the resolution of our data.
@@
where is the value of the model in the wavelength bin and 3.8 is an arbitrary cutoff chosen to fall in the detector gap between NRS1 and NRS2. The prior on metallicity is uniform between 10 and 1000 Solar, and the prior on is uniform between -6 and -1. We do not place a prior on the offset parameters.
For each target we simulate 8 MCMC chains for 1,000,000 steps each, discarding the first 20,000 steps as burn-in (2% of the total simulation). This results in an effective sample size of at least 80,000 for each parameter based on an analysis of the autocorrelation length and parameters equal to 1.0 to three decimal places indicating that the simulations are well-converged, with both quantities computed using arviz (Kumar et al., 2019). We see no evidence of multimodality in the posteriors.
@@
In order to control for the extent to which our methodology contributes to any disagreement between our results and those of previous COMPASS analyses, we also compute the same limits for metallicity and opaque pressure level using identical pipelines from the original works. As described in other works (e.g. Wallack et al., 2024; Teske et al., 2025; Alderson et al., 2025) we compute a similar model grid using PICASO and photochem (Wogan et al., 2025) to compute spectra at chemical equilibrium abundances. We choose two values of C/O (0.5Solar and 1Solar) and 20 log-even spaced metallicities between M/H=[0,3]. We fit for an offset between NRS1 and NRS2 using ultranest (Buchner, 2021). Then, we compute a reduced per data point of the NRS1/NRS2 offset corrected spectrum, which we can then be converted to a p-value and associated -confidence. The effect of clouds is added to the spectrum by including a thick () gray opacity source at a specified pressure level. Our grid includes 5 opaque pressure levels spanning 1-10-4 bars.
IV.3 Composite spectra
@@
V Atmospheric Composition Results
@@
V.1 Influence of C/O Ratio
For planets with equilibrium temperature below about 1000 K, changes in the C/O ratio have minimal impacts on the transmission spectrum (Madhusudhan, 2012). As most of the planets in our sample are in this regime, the results of our thermochemical equilibrium modeling are robust to changes in C/O, justifying the choice to assume C/O=1 Solar. However, TOI-836 b and L 168-9 b, with equilibrium temperatures of 870 K and 1000 K respectively, are warm enough to be influenced by the choice of C/O ratio. At these temperatures the mixing ratios of H2O and CO2 can change substantially between C/O=1 Solar and C/O=1.5 Solar, as illustrated in Figure 18, which in turn drives changes in the posteriors for opaque pressure level versus metallicity.
@@ V.1 Influence of C/O Ratio
In the case of L 168-9 b, for which we find a metallicity lower limit of Solar for the C/O=1 Solar case, the discrepancy is at least in part driven by the planet parameters used in our forward models. Alam et al. (2025) used the mass, radius, and equilibrium temperature from Astudillo-Defru et al. (2020), while we use Hobson et al. (2024). When we substitute the Astudillo-Defru et al. (2020) for our forward model grid computation, we find a metallicity lower limit of 335 Solar for the C/O=1 Solar case. This is significantly closer, but still much higher, than the limit of Solar found by Alam et al. (2025).
@@
While this sensitivity may partially explain the difference between the lower limits derived from the MCMC method and those derived from the -mapping method for L 98-59 c and L 168-9 b (the black and green vertical lines in Figure 20), it cannot explain the discrepancy between the lower limits derived in previous COMPASS studies and those found in this work since the two methods are in such close agreement for the other targets. A possible explanation is found in the fact that the change in the amplitude of transmission features with changes in metallicity grows smaller as metallicity increases. Our L 168-9 b and L 98-59 c spectra constrain the metallicity to be quite high, putting us in a regime where the feature amplitude, and therefore the of the models with respect to the data, are relatively insensitive to changes in metallicity. This means that large changes in the metallicity are necessary to effect small changes in the quality of the fit, making the metallicity lower limit sensitive to small changes in the spectrum between reductions as well as to the goodness of fit metric used (i.e. versus the log-likelihood).
for the metallicity of each COMPASS target for clear atmospheres. The solid red lines represent metallicity lower limits from previous studies of each planet. For TOI-776 b the two red lines represent the reported lower limits for the first and second visits individually, as a lower limit for the combined multi-visit spectrum was not reported in Alderson et al. (2025) The solid black lines correspond to the posterior distribution and lower limit for C/O=1 Solar and the dashed lines are for C/O=1.5 Solar.
V.2 Results of Composite Spectra Analysis
@@
To test this hypothesis we compute two composite spectra: One for the sub-Neptunes TOI-776 c and TOI-836 c, and the other for the five super-Earths (L 98-59 c, L 168-9 b, GJ 357 b, TOI-836 b, and TOI-776 b). Each composite spectra is produced by co-adding the group of spectra with error bars are obtained by summing the error bars for the individual spectra in quadrature. Offsets were not removed prior to co-adding the spectra, so the offset in the composite spectra are the sum of the individual offsets.
@@
Recent work by Kirk and Owen (2025) shows that co-added spectra are unlikely to be informative when the planetary atmospheres in question are insufficiently self-similar. In particular, for equilibrium atmospheres the temperatures should not span too large a range and should not cross any important chemical transitions, such as the transition between atmospheres where CO2 is the dominant Carbon molecule and those atmospheres where CH4 is the dominant Carbon molecule. This occurs at around 600K for planets with Solar C/O ratio. We therefore compute a second set of composite spectra where we have grouped the planets by equilibrium temperature, separating them into a sample of planets with and another with . In this grouping the composite spectrum for the cool planets might be expected to show evidence for a methane feature, while the warmer planets might be expected to show evidence for a CO2 feature. These spectra are shown in Figure 22. We again find no evidence that the spectra are statistically non-flat.
VI Discussion
There are a range of possible explanations for the lack of transmission features for these seven COMPASS targets. The majority of the sample is made up of super-Earths, for which atmosphere detections have proven challenging (see, e.g. May et al., 2023; Cadieux et al., 2024; Moran et al., 2023, as well as previous COMPASS papers). Of the ten super-Earths observed by JWST in transmission in the near-infrared, only one claim of an atmospheric detection has been made, for L 98-59 b (Bello-Arufe et al., 2025). If these planets do possess atmospheres, they may be high mean molecular weight secondary atmospheres. Compositions with mean molecular weights greater than five to ten g/mol are permitted for all five super-Earths in our sample given the precision of the transmission spectra presented here. Future JWST investigations including the Rocky Worlds Director’s Discretionary Time program and the Charting the Cosmic Shoreline Guest Observer program (#7073, PI: Lustig-Yaeger) will help to answer lingering questions about whether these small worlds can harbor atmospheres, at least in the case of the M-dwarf host stars targeted by those programs.
The two sub-Neptunes in our sample, TOI-776 c and TOI-836 c, both have bulk densities that imply the presence of volatiles, and Helium escape has been detected for TOI-836 c, and Hydrogen escape has been observed for TOI-776 demonstrating that they must retain at least some of their initial H/He envelopes (Loyd et al., 2025; Zhang et al., 2025). TOI-836 c is a particularly interesting case due to its proximity in mass-radius space to TOI-421 b, and K2-18 b, both of which have atmosphere detections. However, TOI-836 c has an equilibrium temperature of 665K, which places it in a range where previous studies have suggested that hazes may play a role in attenuating spectral features (Brande et al., 2024; Crossfield and Kreidberg, 2017). Indeed, Wallack et al. (2024) finds that haze production rates on par with those predicted for GJ 1214 b, which is proximate to TOI-836 c in mass, radius, and also temperature, could produce a high-altitude haze for TOI-836 c.
@@ VI Discussion
VI.1 Comparison of COMPASS Targets to other Small Planets Observed with JWST
@@
The seven COMPASS targets studied in this work span from one to three Earth-radii in size. In this same radius range there are a total of six planets observed by JWST in the near-infrared which are thought to show spectral features. These are L 98-59 b (Bello-Arufe et al., 2025), GJ 9827 d (Piaulet-Ghorayeb et al., 2024), TOI-270 d (Benneke et al., 2024), K2-18 b (Madhusudhan et al., 2023), TOI-421 b (Davenport et al., 2025), and GJ 1214 b (Schlawin et al., 2024; Ohno et al., 2025). There are an additional five small planets observed by JWST in the near-infrared which do not have transmission features. These are GJ 3090 b (Ahrer et al., 2025), GJ 1132 b (May et al., 2023; Bennett et al., 2025), GJ 486 b (Moran et al., 2023), LHS 475 b (Lustig-Yaeger et al., 2023), and LHS 1140 b (Cadieux et al., 2024).
@@
In addition to radius, these planets cover similar ranges in mass and equilibrium temperature to the COMPASS targets (see Figures 23 and 24). A possible trend emerges when we consider the distribution of planets with transmission features compared to those without: With a few exceptions (GJ 3090 b and TOI-836 c), the planets with transmission features tend to sit near or above the radius expected for a 50% steam atmosphere (see the dark blue line from Aguichine et al. (2021)) while the planets without transmission features tend to sit below this line, closer to the pure silicate model from Zeng et al. (2016). A fuller understanding of the statistical significance of this trend and its implications for the exoplanet population would require an analysis that is beyond the scope of this work. However, this emerging trend suggests that atmospheres may be challenging to observe for denser worlds.
@@
We examine the transmission spectra that would be produced by the atmospheres of the six planets with features if they were placed around each of the COMPASS planets. We take the retrieved molecular abundances from the original studies, for which best-fit spectra are shown in the left panel of Figure 25, and use these to compute model spectra for the COMPASS targets using petitRADTRANS (Mollière et al., 2019). These models are shown in the right panel of Figure 25, where they are plotted over the observed spectra of each atmosphere. In Figure 26 we show the rejection threshold values between the observed planet spectra (columns) and the model atmosphere (rows). For GJ 1214 b we used the CO2 and CH4 abundances from Schlawin et al. (2024), who did not report a mean molecular weight. We found that in order to reproduce an approximate fit to the observations we had to assume a mean molecular weight of about 42 amu, which is roughly consistent with the reported abundances plus a background gas composed of N2. We note that the composition of GJ 1214 b’s atmosphere and haze properties are highly uncertain (see also Ohno et al. (2025) and Malsky et al. (2025)).
Note that L 98-59 b is not shown in this plot. With a mass and radius of 0.4 and (Demangeon et al., 2021) it sits far below and to the left of the planets shown here.
@@
We find that the low mean molecular weight compositions of K2-18 b and TOI-421 b are inconsistent with the flat spectra of all seven COMPASS targets at greater than 3-. This is because their low mean molecular weight atmospheres produce large transmission features, which are easy to rule out for the COMPASS spectra. TOI-270 d has an intermediate mean molecular weight of =5.82 amu, which results in muted spectral features. Its composition is ruled out for the super-Earths (except for TOI-776 b), but remains barely plausible for TOI-836 c at and is fully consistent with the spectrum of TOI-776 c. The purported high mean molecular weight atmosphere of L 98-59 b is inconsistent with the flat spectrum of its sibling planet L 98-59 c, but is otherwise plausible for the remaining COMPASS targets. However, we note that there are uncertainties about the existence and composition of this atmosphere of L 98-59 b, with the observed JWST spectrum also being consistent with a bare rock (Bello-Arufe et al., 2025). The water-dominated atmosphere of GJ 9827 d is ruled out by the flat spectrum of the super-Earths L 168-9 b and disfavored for TOI-836 b and TOI-776 b, but is otherwise compatible with the remaining COMPASS planets. The very high mean molecular weight, hazy atmosphere of GJ 1214 b is flat enough to be plausible for all of the COMPASS planets.
@@
Summarizing these observations, we see that only the low mean molecular weight atmospheres observed for TOI-421 b and K2-18 b can be reliably ruled out for the entire sample. Higher mean molecular weight atmospheres are largely consistent with the flat spectra observed for these COMPASS targets, with several exceptions in the cases of L 98-59 c and L 168-9 b. It is also worth noting that all of the planets in the left-hand panel of Figure 25 have some short wavelength () coverage, including L 98-59 d and GJ 1214 b which have HST observations that are not shown in the figure. These observations likely help to constrain their atmospheric properties beyond what would be possible with NIRSpec alone, especially due to the lack of strong water features between 2.8 and 5.3 m. In general, the flatness of the COMPASS spectra does not appear all that surprising in light of the wavelength coverage of G395H and the compositional diversity and frequently high mean molecular weight atmospheres already observed among the collection of small planets seen by JWST.
VI.2 Future Observations